Angles x and y are complementary. If x is 20 Degrees more than y, what is the size of angle y?
Answer by checkley77(12844)

X+Y=90 0R X=90-Y
X=Y+20
90-Y=Y+20
-Y-Y=20-90
-2Y=-70
Y=-70/-2
Y=35 DEGREES.
X=35+20=55 DEGREES.
PROOF:
35+55=90
90=90